OpenStudy (anonymous):

Simplify the following expression. 5(x + y) + 3(x - y)

Parth (parthkohli):

Distributive property!

Parth (parthkohli):

Let's just look at \(5(x + y)\).\[5(x+y) =5\cdot x + 5 \cdot y\]

OpenStudy (anonymous):

5x+5y+3x-3y 8x-2y